# Wei Jiang    Jiayu Yang    Yongqi Zhai    Feng Gao    Ronggang Wang

### v9

@@ 3.5 Linear Complexity Intra-Slice Global Context Module
The process is illustrated in Equation 15. In Equation 15, $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$ is employed as the learnable similarity metric, where $10<\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)<1,0<\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1, which makes $10<\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}<$ 1. The non-negativity makes $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$ can be employed as a similarity metric. If $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\to 0$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are near orthogonal. If $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\to 1$, ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ and ${\hat{\boldsymbol{y}}}^{i-1}_{na,q}$ are very similar. The metric is implicit because there is no need to compute $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$. Since we use softmax operation on $\hat{\boldsymbol{y}}^{i-1}_{na,q}$ and $\hat{\boldsymbol{y}}^{i-1}_{ac,k}$ separately, $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\hat{\boldsymbol{y}}^{i}_{ac,v}$ can be computed first during training and testing. The complexity of it is $O(L)$, which is linear with the resolution. The linear complexity makes it easier to employ the global spatial context module for high-resolution image coding.
@@ 3.6 Linear Complexity Inter-Slice Global Context Module
Figure 10: PSNR-Bit-rate curve and MS-SSIM-Bit-rate curve. MS-SSIM is converted to dB for better visual clarity. Please zoom in for better view.
@@
${\boldsymbol{\Phi}}^{i}_{gc,inter}=\textrm{DepthRB}(\hat{\boldsymbol{y}}^{i}_{conv}),$ (18)
where $\hat{\boldsymbol{y}}^{<i}_{q},\hat{\boldsymbol{y}}^{<i}_{k},\hat{\boldsymbol{y}}^{<i}_{v}=\textrm{Embedding}(\hat{\boldsymbol{y}}^{<i})$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2023a) and is employed to enhance the non-linearity. Since the $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{<i}_{k}\right)^{\top}\hat{\boldsymbol{y}}^{<i}_{v}$ is computed during training and testing, the similarity metric is implicit and the overall complexity is $O(L)$, which is linear with the resolution.
Methods Kodak (Kodak, 1993) Tecnick (Asuni & Giachetti, 2014) CLIC Professional Valid (Toderici et al., 2020)
@@
MLIC (ACMMM’23) (Jiang et al., 2023b) $-8.05$ $-49.13$ $-12.73$ $-47.26$ $-8.79$ $-45.79$
${}^{+}$ (ACMMM’23) (Jiang et al., 2023b) $-11.39$ $-52.75$ $-16.38$ $-53.54$ $-12.56$ $-48.75$
${}^{++}$ (Ours) $-13.39$ $-53.63$ $-17.59$ $-53.83$ $-13.08$ $-50.78$
@@
Figure 11: Reconstructions of our proposed ${}^{++}$, recent learned image compression models (He et al., 2022; Zou et al., 2022; Xie et al., 2021; Cheng et al., 2020) and VTM-17.0 Intra (Bross et al., 2021). “[MSE]” denotes the model is optimized for MSE and “[MS-SSIM]” denotes the model is optimized for MS-SSIM (Wang et al., 2003). Please zoom in for better view.
4 Experiments
4.1 Implementation Details
4.1.1 Training Dataset
4.1.2 Training Strategy
${}^{++}$ is built on Pytorch (Paszke et al., 2019) and CompressAI (Bégaint et al., 2020). Following the settings of CompressAI (Bégaint et al., 2020), we set $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE and set $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003). The batch size is set to $32$ and models are trained on a single Tesla A100 GPU. Each model is trained with an Adam optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. We train each model for 2M steps. The learning rate starts at $10^{-4}$ and drops to $3\times 10^{-5}$ at 1.5M steps, drops to $10^{-5}$ at 1.8M steps, and drops to $3\times 10^{-6}$ at 1.9M steps, drops to $10^{-6}$ at 1.95M steps. During training, we random crop images to $256\times 256$ patches during the first $1.2$M steps. To further exploit the effectiveness of global context modules, images are cropped to $512\times 512$ patches during the rest steps. Large patches are beneficial for learning global references. The latent representation can be sparse due to checkerboard partition. A large latent representation makes it more difficult for the model to capture the global contexts and improves the generalization ability of the model across different resolutions.
4.2 Benchmarks and Metrics
  • Kodak (Kodak, 1993) is selected as a test set for almost all end-to-end image compression models (Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Cheng et al., 2020; Minnen & Singh, 2020; Chen et al., 2021; Guo et al., 2022; Wu et al., 2022; Xie et al., 2021; Gao et al., 2021; Chen et al., 2022; Zou et al., 2022; He et al., 2022; Koyuncu et al., 2022; Jiang et al., 2023b; Duan et al., 2023; Liu et al., 2023). It contains 24 raw $768\times 512$ images.
  • Tecnick (Asuni & Giachetti, 2014) contains 100 $1200\times 1200$ RGB images, which is used as a test set in many methods (Ballé et al., 2018; Minnen et al., 2018; Xie et al., 2021; Kim et al., 2022; Jiang et al., 2023b, a; Liu et al., 2023; Zhu et al., 2022b; Duan et al., 2023; Pan et al., 2022).
  • CLIC Professional Valid (Toderici et al., 2020) is the validation set of 3rd Challenge on Learned Image Compression which contains $41$ images. Image in this dataset contains around $2048\times 1440$ pixels. CLIC Pro Valid is widely used in many recent methods (Cheng et al., 2020; Xie et al., 2021; He et al., 2022; Zou et al., 2022; Liu et al., 2023; Jiang et al., 2023b; Wang et al., 2022; Zhu et al., 2022a; Duan et al., 2023).
4.3 Rate-Distortion Performance
4.3.1 Quantitive Results
Rate-distortion curves are presented in Fig. 10. When compared with Cheng’20 (Cheng et al., 2020), our proposed ${}^{++}$ can achieve a maximum improvement of $0.5\sim 0.8$ dB in PSNR and achieve a maximum improvement of $0.6$ dB in MS-SSIM in dB. Our ${}^{++}$ adopts simplified analysis transform and synthesis transform of Cheng’20 (Cheng et al., 2020), therefore, the improvement of model performance is attributed to our linear complexity multi-reference entropy modeling. Our linear complexity multi-reference entropy models can capture more contexts, which leads to much better rate-distortion performance. The improvement also proves correlations exist in multiple dimensions since Cheng’20 (Cheng et al., 2020) adopts an spatial auto-regressive context module. Compared with ELIC (He et al., 2022), our ${}^{++}$ can be up to $0.4$ dB higher at low bit rates in PSNR and $1$ dB higher in MS-SSIM (Wang et al., 2003).
@@ 4.3.1 Quantitive Results
BD-rate reductions are presented in Table 3. When computing BD-rate, VTM-17.0 Intra under YUV444 is employed as anchor. Unofficial weights of ELIC [^2] https://github.com/VincentChandelier/ELiC-ReImplemetation are used to evaluate the rate-distortion performance of ELIC on Tecnick, CLIC Professional Valid. Our ${}^{++}$ outperforms our previous MLIC and ${}^{+}$ (Jiang et al., 2023b). Our ${}^{++}$ reduces BD-rate by $13.39\%$ over VTM-17.0 Intra on Kodak while existing method ELIC only reduces $5.95\%$ BD-rate and LIC-TCM only reduces $10.14\%$ BD-rate on Kodak. Moreover, our ${}^{++}$ performs better on high-resolution datasets, such as Tecnick and CLIC Professional Valid. Our ${}^{++}$ reduces $17.59\%$ on Tecnick and reduces $13.08\%$ BD-rate on CLIC Professional Valid, which are much better than existing methods. Our ${}^{++}$ achieves state-of-the-art performance on all three datasets. The excellent performance of our ${}^{++}$ on these datasets also demonstrates the excellent generalization of our ${}^{++}$. We highlight the BD-rate for MS-SSIM, our proposed ${}^{++}$ reduces about $50\%$ bits compared to VTM-17.0 Intra, which is a large progress in learned image compression.
@@ 4.3.2 Qualitive Results
To further demonstrate superiority of our proposed ${}^{++}$, we compare our ${}^{++}$ with learned image compression models Cheng’20 (Cheng et al., 2020), Xie’21 (Xie et al., 2021), STF (Zou et al., 2022), WACNN (Zou et al., 2022), ELIC (He et al., 2022) and non-neural codec VTM-17.0 Intra (Bross et al., 2021) on perceptual quality. Fig. 11 presents the reconstructions of Kodim07 from Kodak. PSNR value of the image reconstructed by our ${}^{++}$ optimized for MSE is $1$dB higher than image reconstructed by VTM-17.0 Intra (Bross et al., 2021). MS-SSIM of the image reconstructed by our ${}^{++}$ optimized for MS-SSIM is $0.02$ higher than image reconstructed by VTM-17.0 Intra. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of ${}^{++}$ have sharper textures retain more details. In terms of visual quality, our ${}^{++}$ have significant improvements on rate-perception performance compared to other models.
Figure 12: GPU memory consumption, encoding time, decoding time comparisons among our proposed models (Cheng et al., 2020; Xie et al., 2021; Qian et al., 2022; Zou et al., 2022; He et al., 2022; Liu et al., 2023) and recent learned image compression models. Encoding and decoding time of Qian’21 (Qian et al., 2020), Entroformer (Qian et al., 2022), NeuralSyntax (Wang et al., 2022) exceed 4.5s.
Context Modules ${}^{++}$ Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8
@@
Table 4: Ablation Studies on Kodak. BD-rate (%) is employed to evaluate their contributions. ${}^{++}$ is the anchor.
@@ 4.4 Computational Complexity
In order to better fit the real-world scenario, we test the model complexities in the case of different resolution images as inputs. We select $16$ images with resolution larger than $3584\times 3584$ from LIU4K test dataset (Liu et al., 2020a) and we center crop these images to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072,3584\times 3584\}$ patches. We compared our ${}^{++}$ with our prior work ${}^{+}$ which employs vanilla attention to illustrate the advantages of proposed linear complexity global context capturing. We also compare our proposed ${}^{++}$ with recent learned image compression models (Cheng et al., 2020; Qian et al., 2020; Xie et al., 2021; Qian et al., 2022; Wang et al., 2022; Zou et al., 2022; He et al., 2022; Liu et al., 2023). The experiments are conducted on a Tesla A100 GPU and a Xeon(R) Platinum 8260C CPU. The results are illustrated in Fig. 12.
@@ 4.4.1 On GPU Memory comsuption
The quadratic complexity of vanilla attention leads to significantly more memory consumptions on high resolution image coding. When compressing $2048\times 2048$ images, ${}^{+}$ consumes nearly $22.37$ GB GPU memory. Our proposed linear complexity global context capturing significantly reduces the consumption of GPU memory as our proposed ${}^{++}$ only takes about $5.55$ GB GPU memory to compress a $2048\times 2048$ image. When compressing $3072\times 3072$ images, ${}^{+}$ consumes $46.38$ GB GPU memory while our proposed ${}^{++}$ only consumes $7.69$ GB GPU memory. Compared with recent LIC-TCM (Liu et al., 2023), our ${}^{++}$ consumes $\frac{1}{2}$ of the GPU memory consumed by LIC-TCM when compressing $2048\times 2048$ images. When compressing $2560\times 2560$ images, our ${}^{++}$ only consumes about $\frac{1}{4}$ of the GPU memory consumed by LIC-TCM. When compressing a $3584\times 3584$ image, the GPU memory consumption of our proposed ${}^{++}$ is still only $12.3$ GB while the LIC-TCM requires $45.95$ GB GPU memory. The curve of the GPU memory consumed by our ${}^{++}$ as the resolution grows is also much flatter.

### v10

@@ 3.5 Linear Complexity Intra-Slice Global Context Module
The metric is implicit because there is no need to compute $\textrm{softmax}_{2}\left(\hat{\boldsymbol{y}}^{i-1}_{na,q}\right)\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}$. Since we use softmax operation on $\hat{\boldsymbol{y}}^{i-1}_{na,q}$ and $\hat{\boldsymbol{y}}^{i-1}_{ac,k}$ separately, $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{i-1}_{ac,k}\right)^{\top}\hat{\boldsymbol{y}}^{i}_{ac,v}$ can be computed first during training and testing. The complexity of it is $O(L)$, which is linear with the resolution. The linear complexity makes it easier to employ the global spatial context module for high-resolution image coding.
@@ 3.6 Linear Complexity Inter-Slice Global Context Module
$\hat{\boldsymbol{y}}^{i}_{conv}=\textrm{conv}_{K\times K}(\hat{\boldsymbol{y}}^{i}_{attn}),$
${\boldsymbol{\Phi}}^{i}_{gc,inter}=\textrm{DepthRB}(\hat{\boldsymbol{y}}^{i}_{conv}),$
where $\hat{\boldsymbol{y}}^{<i}_{q},\hat{\boldsymbol{y}}^{<i}_{k},\hat{\boldsymbol{y}}^{<i}_{v}=\textrm{Embedding}(\hat{\boldsymbol{y}}^{<i})$, Embedding is the embedding layer. Embedding layer consists of a $1\times 1$ convolutional layer and a $3\times 3$ depth-wise convolutional layer. The $3\times 3$ depth-wise convolutional layer is employed for learnable position embedding. DepthRB is the depth-wise residual bottleneck (Jiang et al., 2024) and is employed to enhance the non-linearity. Since the $\textrm{softmax}_{1}\left(\hat{\boldsymbol{y}}^{<i}_{k}\right)^{\top}\hat{\boldsymbol{y}}^{<i}_{v}$ is computed during training and testing, the similarity metric is implicit and the overall complexity is $O(L)$, which is linear with the resolution.
4 Experiments
4.1 Implementation Details
4.1.1 Training Dataset
4.1.2 Training Strategy
MLIC++ is built on Pytorch (Paszke et al., 2019) and CompressAI (Bégaint et al., 2020). Following the settings of CompressAI (Bégaint et al., 2020), we set $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE and set $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003). The batch size is set to $32$ and models are trained on a single Tesla A100 GPU. Each model is trained with an Adam optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$. We train each model for 2M steps.
Figure 10: PSNR-Bit-rate and MS-SSIM-Bit-rate curves. MS-SSIM is converted to dB for better clarity. Please zoom in for better view.
The learning rate starts at $10^{-4}$ and drops to $3\times 10^{-5}$ at 1.5M steps, drops to $10^{-5}$ at 1.8M steps, and drops to $3\times 10^{-6}$ at 1.9M steps, drops to $10^{-6}$ at 1.95M steps. During training, we random crop images to $256\times 256$ patches during the first $1.2$M steps. To further exploit the effectiveness of global context modules, images are cropped to $512\times 512$ patches during the rest steps. Large patches are beneficial for learning global references. The latent representation can be sparse due to checkerboard partition. A large latent representation makes it more difficult for the model to capture the global contexts and improves the generalization ability of the model across different resolutions.
4.2 Benchmarks and Metrics
  • Kodak (Kodak, 1993) is selected as a test set for almost all end-to-end image compression models (Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Cheng et al., 2020; Minnen & Singh, 2020; Chen et al., 2021; Guo et al., 2022; Wu et al., 2022; Xie et al., 2021; Gao et al., 2021; Chen et al., 2022; Zou et al., 2022; He et al., 2022; Koyuncu et al., 2022; Jiang et al., 2023; Duan et al., 2023b; Liu et al., 2023). It contains 24 raw $768\times 512$ images.
  • Tecnick (Asuni & Giachetti, 2014) contains 100 $1200\times 1200$ RGB images, which is used as a test set in many methods (Ballé et al., 2018; Minnen et al., 2018; Xie et al., 2021; Kim et al., 2022; Jiang et al., 2023, 2024; Liu et al., 2023; Zhu et al., 2022b; Duan et al., 2023b; Pan et al., 2022).
  • CLIC Professional Valid (Toderici et al., 2020) is the validation set of 3rd Challenge on Learned Image Compression which contains $41$ images. Image in this dataset contains around $2048\times 1440$ pixels. CLIC Pro Valid is widely used in many recent methods (Cheng et al., 2020; Xie et al., 2021; He et al., 2022; Zou et al., 2022; Liu et al., 2023; Jiang et al., 2023; Wang et al., 2022; Zhu et al., 2022a; Duan et al., 2023b).
4.3 Rate-Distortion Performance
4.3.1 Quantitive Results
Rate-distortion curves are presented in Fig. 10. When compared with Cheng’20 (Cheng et al., 2020), our proposed MLIC++ can achieve a maximum improvement of $0.5\sim 0.8$ dB in PSNR and achieve a maximum improvement of $0.6$ dB in MS-SSIM in dB. Our MLIC++ adopts simplified analysis transform and synthesis transform of Cheng’20 (Cheng et al., 2020), therefore, the improvement of model performance is attributed to our linear complexity multi-reference entropy modeling. Our linear complexity multi-reference entropy models can capture more contexts, which leads to much better rate-distortion performance. The improvement also proves correlations exist in multiple dimensions since Cheng’20 (Cheng et al., 2020) adopts an spatial auto-regressive context module. Compared with ELIC (He et al., 2022), our MLIC++ can be up to $0.4$ dB higher at low bit rates in PSNR and $1$ dB higher in MS-SSIM (Wang et al., 2003).
BD-rate reductions are presented in Table 3.
Methods Kodak Tecnick CLIC Professional Valid
@@ 4.3.1 Quantitive Results
MLIC (ACMMM’23) (Jiang et al., 2023) $-8.05$ $-49.13$ $-12.73$ $-47.26$ $-8.79$ $-45.79$
MLIC+ (ACMMM’23) (Jiang et al., 2023) $-11.39$ $-52.75$ $-16.38$ $-53.54$ $-12.56$ $-48.75$
QARV (TPAMI’24) (Duan et al., 2023a) $+0.31$- $-3.03$- -- --
FTIC (ICLR’24) (Li et al., 2024) $-12.99$ $-51.13$ $-14.88$- $-9.53$-
LLIC-TCM (TMM’24) (Jiang et al., 2024) $-10.94$ $-49.73$ $-14.99$- $-10.41$-
MLIC++ (Ours) $-13.39$ $-53.63$ $-17.59$ $-53.83$ $-13.08$ $-50.78$
@@
Figure 11: Reconstructions of our proposed MLIC++, recent learned image compression models (He et al., 2022; Zou et al., 2022; Xie et al., 2021; Cheng et al., 2020) and VTM-17.0 Intra (Bross et al., 2021). “[MSE]” denotes the model is optimized for MSE and “[MS-SSIM]” denotes the model is optimized for MS-SSIM (Wang et al., 2003). Please zoom in for better view.
@@
When computing BD-rate, VTM-17.0 Intra under YUV444 is employed as anchor. Unofficial weights of ELIC [^2] https://github.com/VincentChandelier/ELiC-ReImplemetation are used to evaluate the rate-distortion performance of ELIC on Tecnick, CLIC Professional Valid. Our MLIC++ outperforms our previous MLIC and MLIC+ (Jiang et al., 2023). Our MLIC++ reduces BD-rate by $13.39\%$ over VTM-17.0 Intra on Kodak while existing method ELIC only reduces $5.95\%$ BD-rate and LIC-TCM only reduces $10.14\%$ BD-rate on Kodak. Moreover, our MLIC++ performs better on high-resolution datasets, such as Tecnick and CLIC Professional Valid. Our MLIC++ reduces $17.59\%$ on Tecnick and reduces $13.08\%$ BD-rate on CLIC Professional Valid, which are much better than existing methods. Our MLIC++ achieves state-of-the-art performance on all three datasets. The excellent performance of our MLIC++ on these datasets also demonstrates the excellent generalization of our MLIC++. We highlight the BD-rate for MS-SSIM, our proposed MLIC++ reduces about $50\%$ bits compared to VTM-17.0 Intra, which is a large progress in learned image compression.
@@ 4.3.2 Qualitive Results
To further demonstrate superiority of our proposed MLIC++, we compare our MLIC++ with learned image compression models Cheng’20 (Cheng et al., 2020), Xie’21 (Xie et al., 2021), STF (Zou et al., 2022), WACNN (Zou et al., 2022), ELIC (He et al., 2022) and non-neural codec VTM-17.0 Intra (Bross et al., 2021) on perceptual quality. Fig. 11 presents the reconstructions of Kodim07 from Kodak. PSNR value of the image reconstructed by our MLIC++ optimized for MSE is $1$dB higher than image reconstructed by VTM-17.0 Intra (Bross et al., 2021). MS-SSIM of the image reconstructed by our MLIC++ optimized for MS-SSIM is $0.02$ higher than image reconstructed by VTM-17.0 Intra. The windowsill of reconstructions are cropped to patches for clearer comparisons. Upon closer inspection, certain regions within the images show more pronounced differences. For instance, in the red-circled areas of Figure 11. Moreover, our MSE-optimized MLIC++ model achieves lower bpp values—ranging from $84.4\%$ to $93.8\%$ of those in other models like ELIC, STF, WACNN, and Cheng’20—while maintaining competitive image quality. This demonstrates the efficiency and effectiveness of our approach.
Figure 12: GPU memory consumption, encoding time, decoding time comparisons among our proposed models and recent learned image compression models (Cheng et al., 2020; Xie et al., 2021; Qian et al., 2022; Zou et al., 2022; He et al., 2022; Liu et al., 2023). Encoding and decoding time of Qian’21 (Qian et al., 2020), Entroformer (Qian et al., 2022), NeuralSyntax (Wang et al., 2022) exceed 4.5s. Please zoom in for better view.
Context Modules MLIC++ Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8
@@
Table 4: Ablation Studies on Kodak. BD-rate (%) is employed to evaluate their contributions. MLIC++ is the anchor.
@@ 4.4 Computational Complexity
In order to better fit the real-world scenario, we test the model complexities in the case of different resolution images as inputs. We select $16$ images with resolution larger than $3584\times 3584$ from LIU4K test dataset (Liu et al., 2020a) and we center crop these images to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072,3584\times 3584\}$ patches. We compared our MLIC++ with our prior work MLIC+ which employs vanilla attention to illustrate the advantages of proposed linear complexity global context capturing. We also compare our proposed MLIC++ with recent learned image compression models (Cheng et al., 2020; Qian et al., 2020; Xie et al., 2021; Qian et al., 2022; Wang et al., 2022; Zou et al., 2022; He et al., 2022; Liu et al., 2023). The experiments are conducted on a Tesla A100 GPU and a Xeon(R) Platinum 8260C CPU. The results are illustrated in Fig. 12.
@@ 4.4.1 On GPU Memory comsuption
The quadratic complexity of vanilla attention leads to significantly more memory consumptions on high resolution image coding. When compressing $2048\times 2048$ images, MLIC+ consumes nearly $22.37$ GB GPU memory. Our proposed linear complexity global context capturing significantly reduces the consumption of GPU memory as our proposed MLIC++ only takes about $5.55$ GB GPU memory to compress a $2048\times 2048$ image. When compressing $3072\times 3072$ images, MLIC+ consumes $46.38$ GB GPU memory while our proposed MLIC++ only consumes $7.69$ GB GPU memory. Compared with recent LIC-TCM (Liu et al., 2023), our MLIC++ consumes $\frac{1}{2}$ of the GPU memory consumed by LIC-TCM when compressing $2048\times 2048$ images. When compressing $2560\times 2560$ images, our MLIC++ only consumes about $\frac{1}{4}$ of the GPU memory consumed by LIC-TCM. When compressing a $3584\times 3584$ image, the GPU memory consumption of our proposed MLIC++ is still only $12.3$ GB while the LIC-TCM requires $45.95$ GB GPU memory. The curve of the GPU memory consumed by our MLIC++ as the resolution grows is also much flatter.